Find the equation of the hyperbola with vertices (1,−4) and (1,4), and foci at(1,−6) and (1,6).
1
Expert's answer
2020-05-18T16:08:39-0400

One thing to notice in all the coordinates is that their x coordinate are same i.e. all the points lie on a line x=1

hence the center of the hyperbola = (1,(-4+4)/2) = (1,0)

distance between center and vertex (a) = 4

the foci are at distance 6 from the center , hence c=6


"a^2+b^2=c^2 \\b^2=20"


"\frac{(y-k)^2}{a^2}-\frac{(x-h)^2}{b^2}=1"


"\frac{(y-0)^2}{16}-\frac{(x-1)^2}{20}=1"